# Logical Phase Transitions: Understanding Collapse in LLM Logical Reasoning

### v1

@@ B.3 Example of NSA-LR Data Instance
B.4 Baselines
Here we illustrate the details of each baseline used for comparison. The full prompts are in Section F.
Naive Prompting
This approach involves directly presenting the question to the model without any additional instructions or intermediate reasoning steps. The model is expected to produce the final answer based solely on its pretrained knowledge. It does not encourage or structure the reasoning process, making it suitable only for simple or factual queries.
Chain-of-Thought (CoT)
CoT prompting introduces intermediate reasoning steps by encouraging the model to articulate its thought process before reaching a conclusion. Instead of predicting the final answer directly, the prompt guides the model to perform step-by-step reasoning. This has been shown to improve multi-step reasoning performance on complex tasks (COT-2022chain).
@@ Appendix C Experimental Details
Overview. This section provides detailed experimental settings and extended analyses to ensure reproducibility and support the main findings. We describe the evaluation protocol and implementation details (§C.1), report full results across Model $\times$ Complexity BinsC.2), and present diagnostic analyses of the complexity metric and logical phase transitions (§C.5, §C.4). We further examine potential confounding factors, including complexity–premise count relations (§C.6) and reasoning effort (§C.7).
@@ C.2 Full Results
C.3 Necessity of Curriculum Training
To examine the necessity of curriculum learning across logical complexity regimes, we conduct controlled fine-tuning experiments in which training is restricted to a single complexity interval. As shown in Figure 7, models trained exclusively on high-complexity instances (hard-only) exhibit improved accuracy within the high-complexity region, but experience a substantial performance degradation on low- and medium-complexity samples. This behavior indicates catastrophic forgetting of reasoning patterns acquired in simpler regimes.
In contrast, training restricted to low- or medium-complexity instances does not improve performance on high-complexity samples and in some cases leads to further degradation in that regime. These results show that training on a single complexity regime is insufficient to maintain consistent performance across the full complexity spectrum, motivating the use of complexity-aware curriculum training.
@@ C.4 Logical Phase Transition Analysis
C.5 LoCM Diagnostics and Calibration
@@
C.6 LoCM vs. Premise Count
@@
C.7 Reasoning Effort Analysis
@@ D.2 Well-Formedness and CFG Validation
CFG grammar. To ensure the well-formedness of all FOL expressions, we implement a symbolic parser using nltk (tool-nltk-2006nltk) and validate formulas with a context-free grammar (CFG):
@@ Appendix E Future Directions
LoCM as a diagnostic tool for reasoning robustness. Building on the view of Russell Principles-mathematics2020russell and Wittgenstein logico-philosophicus2023wittgenstein that logic underlies the structure of the world, we treat LoCM not merely as a metric for symbolic tasks but as a general framework for measuring reasoning complexity across domains. Any domain with an explicit logical structure can be formalized into symbolic reasoning chains. By encoding domain-specific premises and inference rules in an LoCM-compatible form, researchers can identify model-specific complexity thresholds, positioning LoCM as a diagnostic tool for reasoning robustness and for analyzing how LLMs engage with the logical structure of knowledge.
@@ Appendix F Full Prompts and Examples
Below are detailed prompts used in our evaluation.The prompts for constructing our datasets can be found in ProverGen (Benchmark-ProverQA-ICLR-2025large).

### v2

@@ Appendix C Experimental Details
Overview. This section provides detailed experimental settings and extended analyses to ensure reproducibility and support the main findings. We describe the evaluation protocol and implementation details (§C.1), report full results across Model $\times$ Complexity BinsC.2), analyze the necessity of curriculum design (§C.3), include an analysis of the FOLIO CoT anomaly (§C.4), and present diagnostic analyses of the complexity metric and logical phase transitions (§C.5, §C.6). We further examine potential confounding factors, including complexity–premise count relations (§C.7) and reasoning effort (§C.8).
@@ C.2 Full Results
C.3 Necessity of Curriculum Training
Single-regime training is insufficient for robust reasoning across logical complexity levels. To examine this, we conduct controlled fine-tuning experiments in which training is restricted to a single complexity interval. As shown in Figure 7, models trained exclusively on high-complexity instances (hard-only) achieve better accuracy in the high-complexity region but suffer substantial degradation on low- and medium-complexity samples, indicating catastrophic forgetting of reasoning patterns associated with simpler regimes. By contrast, training restricted to low- or medium-complexity instances does not improve performance on high-complexity samples and can even further degrade that regime. Taken together, these results show that single-regime training cannot sustain consistent reasoning performance across complexity levels, motivating the use of complexity-aware curriculum training.
@@ C.3 Necessity of Curriculum Training
C.4 Analysis of the FOLIO CoT Anomaly
As shown in Table 2, FOLIO is the main exception to the otherwise consistent gains of NSCT under CoT prompting. We do not view this as evidence of instability specific to NSCT, since similar drops on FOLIO also appear for several models fine-tuned on formal reasoning datasets, including the ProntoQA-tuned, ProofWriter-tuned, and FOLIO-tuned variants. This suggests that the anomaly more likely reflects a broader mismatch between formal-logic-oriented tuning and the reasoning demands of FOLIO.
Although FOLIO is annotated with first-order logic, it places heavier demands on natural-language interpretation and commonsense grounding than the more structurally controlled benchmarks. In contrast, NSCT is primarily designed to strengthen formal validity under increasing logical complexity. We therefore interpret the FOLIO result as a boundary case of the current framework rather than a contradiction of the main findings, and view it as a motivation for incorporating more commonsense-sensitive reasoning data in future work.
C.5 Logical Phase Transition Analysis
@@
C.6 LoCM Diagnostics and Calibration
@@
C.7 LoCM vs. Premise Count
@@
C.8 Reasoning Effort Analysis
@@ D.2 Well-Formedness and CFG Validation
CFG grammar. To ensure the well-formedness of all FOL expressions, we implement a symbolic parser using nltk (Bird, 2006) and validate formulas with a context-free grammar (CFG):
@@ Appendix E Future Directions
LoCM as a diagnostic tool for reasoning robustness. Building on the view of Russell Russell (2020) and Wittgenstein Wittgenstein (2023) that logic underlies the structure of the world, we treat LoCM not merely as a metric for symbolic tasks but as a general framework for measuring reasoning complexity across domains. Any domain with an explicit logical structure can be formalized into symbolic reasoning chains. By encoding domain-specific premises and inference rules in an LoCM-compatible form, researchers can identify model-specific complexity thresholds, positioning LoCM as a diagnostic tool for reasoning robustness and for analyzing how LLMs engage with the logical structure of knowledge.
@@ Appendix F Full Prompts and Examples
Below are detailed prompts used in our evaluation.The prompts for constructing our datasets can be found in ProverGen (Qi et al., 2025).